OpenStudy (anonymous):

Nestor borrows $10,000 from the shady loan shark, Mr. Wildcat. Mr. Wildcat tells him he must repay the loan in 1 month and pay an additional $1000 in interest. If Mr. Wildcat is charging simple interest on his loan, what is the annual interest rate?

OpenStudy (anonymous):

@Calcmathlete help? :)

OpenStudy (anonymous):

r=I/Pt 1000/(10,000)(1)=0.1 is that right or wrong? (any of it lol)

OpenStudy (anonymous):

I = Prt 1000 = 10000(x)(1/12) THis answer is the decimal value, so multiply it by 100 to get the percent value.

OpenStudy (anonymous):

1.2000048 thats what i got?

OpenStudy (anonymous):

It's just 1.2. Still, I believe that you got it.

OpenStudy (anonymous):

do i multiply 1.2 by 100 now?

OpenStudy (anonymous):

If you want the percentage value, yes.

OpenStudy (anonymous):

120%? that doesn't make sense to me...

OpenStudy (anonymous):

That's what I got, but I believe it works because to pay the lone is 1 month, the time part of the formula is 1/12.

OpenStudy (anonymous):

okay thanks.

OpenStudy (anonymous):

np :)